OpenStudy (anonymous):

Let f be a linear function where f(2)=5 and f(-3)=1 State the function f(x)

OpenStudy (anonymous):

I have a function for the first one. f(x)=3+x f(2)=5? Is that a good one?

OpenStudy (anonymous):

linear functions are of the form: f(x) = ax+b so if f(2) = 5 means when you plug in 2 to x you get 5, then 2*a+b=5, similarly -3*a+b=1 now solve these: 2*a+b=5 => a=(5-b)/2 and plug it into -3*[(5-b)/2]+b=1 and solve for b,

OpenStudy (anonymous):

or you can solve thee two equations simultanesouly and eliminate one of the unknowns, say a, then you can plug the found b value into one of the equations to get a

OpenStudy (anonymous):

so mine wouldn't work would it be correct to use two separate equations?

OpenStudy (anonymous):

or one single equation

OpenStudy (anonymous):



OpenStudy (anonymous):

i didnt get the way you found f(x) = 3+x

OpenStudy (anonymous):

well here's the thing say we have f(x)=3+x we plug in f(2)= 3+2 which =5 therefore f(2)=5

OpenStudy (anonymous):

but that does not work for f(-3) = 1, if the function were f(x)=3+x then we would expect f(-3) = 3 + (-3) to be 0.

OpenStudy (anonymous):

ahh i understand now.

OpenStudy (anonymous):

the function should work for any values so the point is: if a function is "linear", just write it of the form ax+b and plug the given values of x and make them equal to the results given.. you'll then have two different equations with a and b in them (because we already replaced xs with the values given) and solve them simultnesouly, find a and b seperately and construct the function such that f(x) = a*x + b

OpenStudy (anonymous):

i have a question in your work why is there a negative in the line 2a+b=5

OpenStudy (xapproachesinfinity):

Do you know about slope, you can use that to find the equation of that line. Much easier and a shortcut to the solution

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok, when you write two different equations that way, you should ask yourself which of the unknowns i can eliminate; the easiest way for those equations was to multiply the first one by (-1) and add it to the one below, then we got rid of b's! we found a and plugged the a values to find b

OpenStudy (anonymous):

okay i got it soo would this be the equation?

OpenStudy (anonymous):